OpenStudy (lachegapa):

Pin Numbers A student loan administrator distributes pin numbers to its debtors. Each pin consists of two letters followed by three numbers. (Assume repetition is allowed, and order is important.) (a) How many different pin numbers are there? (b) What is the probability a pin number selected at random ends in 000? (c) What is the probability a pin number selected at random begins with AA?

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (lachegapa):

Thank you.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

counting principle for this one how many letters in the alphabet?

OpenStudy (mhchen):

(A) 2 letter and 3 numbers, repetition is allowed (26 letters) * (26 letters) * (10 numbers) * (10 numbers) * (10 numbers) = 26 * 26 * 10 * 10 * 10 (B) Number of ways to end in 000: (26 letters) * (26 letters) * (1 number) * (1 number) * (1 number) = 26 * 26 * 1 * 1 * 1 Total = (A) = 26 * 26 * 10 * 10 * 10 Probability: (26*26*1*1*1)/(26*26*10*10*10) (C)Number of ways to being with AA: (1 letter) * (1 letter) * (26 letters) * (10 numbers) * (10 numbers) Now I think you know how to do it :#